The capital structure of ABC Pvt. Ltd is as follows:

Equity share capital (eachshareofRs.10) = Rs.16, 00,000

Debentures with a coupon rate of 10% = Rs. 10, 00,000

Reserves and surplus = Rs.15, 00,000

Revenue from the business activities for the company is Rs. 2.00 crores. Its variable cost is 10% of the revenue, fixed operating cost is Rs. 60 lakhs and the company pays income tax at a rate of 25%. (10Marks)

a. Calculate financial leverage, operating leverage and combined leverage for the company.

b. Determine the likely level of EBIT for EPS of (i) Rs.45, (ii) Rs.60, and(iii) Rs. 75.


1
Expert's answer
2021-08-30T14:32:59-0400

Solution:

a.). Financial Leverage ="\frac{Total \; Debt}{Shareholders\; Equity}"

Total Debt = 1000,000

Shareholders Equity = Equity shares + Reserves and Surplus = 1,600,000 + 1,500,000 = 3,100,000


Financial Leverage = "\frac{1,000,000}{3,100,000} = 0.32"


Financial Leverage = 0.32

 

Operating Leverage = "\frac{Fixed \; Costs}{Total\; Costs}"


Fixed Costs = 60 lakhs = 0.60 crores

Variable Costs = "10\%\times 2" crores = 0.20 crores

Total Costs = FC + VC = 0.60 + 0.20 = 0.80 crores


Operating Leverage = "\frac{0.60}{0.80} = 0.75"


Operating Leverage = 0.75

 

Combined Leverage = Operating Leverage x Financial Leverage = 0.75 "\times" 0.32 = 0.24

 

Combined Leverage = 0.24


b.). EBIT for EPS of:

i.). Rs.45

EPS = "\frac{Net \; Income}{No\;of \;Shares\;Outstanding}"


No. of shares outstanding = "\frac{1,600,000}{10}" = 160,000 shares


45 = "\frac{Net \; Income}{160,000}"

Net income = 45 x 160,000                = 7,200,000

Less Income Tax (25"\% \times" 7,200,000) = (1,800,000)

Less Interest (10"\% \times"1,000,000)       =   (100,000)

EBIT                                                   = 5,300,000

 

ii.). Rs.60

60 = "\frac{Net \; Income}{160,000}"

Net income = 60 "\times" 160,000                = 9,600,000

Less Income Tax ("25\% \times"7,200,000) = (2,400,000)

Less Interest ("10\% \times" 1,000,000)       =   (100,000)

EBIT                                                   = 7,100,000

 

ii.). Rs.75

75 = "\frac{Net \; Income}{160,000}"

Net income = 75 "\times"160,000          = 12,000,000

Less Income Tax ("25\% \times" 25% x 7,200,000) = (3,000,000)

Less Interest (10"\% \times" 1,000,000)        =   (100,000)

EBIT                                                   = 8,900,000